# Fixing non-blocking data structures for better compatibility with memory reclamation schemes

### v1

@@ Abstract
We present a new technique, Safe Concurrent Optimistic Traversals (SCOT), to address a well-known problem with Hazard Pointers (HP) related to optimistic traversals. Unlike Epoch-Based Reclamation (EBR), HP enables strong robustness properties but lacks support for some data structures such as Harris’ original list and Natarajan-Mittal tree, which can slow down operations substantially (e.g., when replacing with Harris-Michael list) or make correct implementations impossible altogether (e.g., Natarajan-Mittal tree).
Recent works propose replacing HP with another robust memory reclamation schemes that support optimistic traversals such as HP++. However, HP++ is generally slower than HP, has undesirable applicability trade-offs, and more complex implementation, among other issues. We propose a different method which keeps HP intact but instead relies on data structure adaptations. Moreover, our method applies to other recent (and faster) robust reclamation schemes such as Hazard Eras (HE) and Interval-Based Reclamation (IBR), which are HP-like and have similar problems related to optimistic traversals.
We implement two fundamentally different data structures: Harris’ list and Natarajan-Mittal tree. Our method enables their first correct implementations with optimistic traversals with HP and IBR. Our method is also directly applicable to other data structures which extend Harris’ list, e.g., skip lists and hash maps. We observe similar performance benefits of Harris’ original list (vs. Harris-Michael list) that were previously only available to EBR users. As we discuss in the paper, our Natarajan-Mittal tree is the first bug-free implementation for HP and IBR; it achieves very high throughput, comparable to that of EBR.
@@ 1 Introduction
When using manual memory management, memory has to be safely reclaimed by freeing memory only when all ongoing operations with stale pointers complete. A number of past researchers proposed various safe memory reclamation (SMR) techniques in C, C++, or Rust. Non-blocking SMR progress is often crucial for the overall progress guarantees, as we do not want to compromise data structure’s progress properties. Unfortunately, many SMR schemes are trying to attain several highly desirable properties simultaneously while often unknowingly compromising non-blocking progress or other important properties one way or the other. A recent work [21] sheds more light onto this conundrum: it proves that it is impossible to achieve three highly desirable SMR properties simultaneously, i.e., only two properties can ever be achieved: (A) robustness (bounded memory usage), (B) easy integration (no dependency on any special APIs, such as roll-back mechanisms or OS-based primitives), and (C) wide applicability (compatibility with most non-blocking data structures).
@@
While (A) is non-negotiable for non-blocking progress, [^1] If memory usage is unbounded, no further system progress can be made once memory is exhausted, regardless of whether actual locks are present. past researchers would compromise either (B), (C), or other properties that are tangentially related to non-blocking progress, sometimes without fully realizing that. To make matters worse, many researchers prioritize (C) because, intuitively, their goal is to support as many data structures as possible. However, we argue that sacrificing (C) rather than (B) by carefully adapting existing problematic data structures is a “necessary evil”, which is a crucial new insight. First, (C) in its strongest form – applying to all data structures – only currently known to exist in blocking SMR schemes [21], which is a non-starter for us. Second, even if we accept (C) in its relaxed form – applying to most data structures – the absence of (B) would lead to either inability to achieve non-blocking progress in practice due to dependence on OS-based primitives, lack of universally-defined APIs, or both. (We discuss some specific examples in Section 5.) Third, the absence of (C) simply implies that adaptations are required, not impossibility of the implementation.
@@
EBR [9] is a fast and easy-to-use approach but has a serious drawback that any stalled thread causes unbounded memory usage, i.e., it lacks (A). Once exhausting memory, the system must either reboot or block until memory is freed. Many researchers prefer EBR’s simplicity but attempt to solve its robustness issue. Unfortunately, such approaches typically sacrifice strict non-blocking progress [5, 22, 4, 1, 2], as discussed in Section 5, or need non-trivial roll-back mechanisms [20], which are fairly hard to generalize in terms of the API within the same library framework.
@@
HP [14] is an antipode of EBR: it is somewhat slower and more difficult to use but provides robustness with strict non-blocking guarantees by placing a hard limit on the number of local pointers within each thread. Prior to this work, well-known data structures such as Harris’ linked list [10] with optimistic traversals [^2] The original paper [10] did not explicitly mention that search can be done in a read-only manner (w/o any CAS operations). In this paper, by Harris’ list, we denote the enhanced version [11] with optimistic read-only traversals. and Natarajan-Mittal’s binary search tree (BST) [15] were not known to work [^3] Despite presenting Natarajan-Mittal Tree results in [25, 17] for HP and other SMR schemes, it was later argued [3] that such implementations are buggy and leak memory. Moreover, [3] pointed out two bugs, and only one bug was resolved in [3]. Prior to our work, the second bug related to optimistic traversals remained unresolved. with HP [12, 3], which is due to HP’s lack of the wide-applicability (C) property. A number of other recent schemes, including Interval Based Reclamation (IBR) and Hazard Eras (HE), sometimes offer better performance but have similar issues to HP. HP++ [12] is another recent scheme. HP++ is slower than HP and, strictly speaking, makes compromises with both (B) and (C). However, HP++ still demonstrates support for these two data structures. We are inspired by HP++’s success with these data structures, which is indicative that many existing data structures can still be implemented, with an additional effort, despite prior beliefs regarding limited applicability of the SMR schemes in this category.
@@
In this paper, we depart from a typical strategy of designing a “silver-bullet” SMR approach for the reasons described above. Instead, we argue that the existing problematic data structures can be modified to accommodate SMR schemes with strict non-blocking properties. We demonstrate our new method in Section 4 for Harris’ list, Natarajan-Mittal tree, and other data structures. In Section 5, we also touch upon how limitations of our approach with respect to wait-free traversals can be addressed.
@@ 2.1 Lock- and Wait-Freedom
The most widely used non-blocking progress is lock-freedom, which allows threads to interfere with each other while still guaranteeing that at least one thread makes progress in a finite number of steps. This should not be confused with “lockless” approaches that simply avoid explicit locks, where one preempted thread can block all other threads. Wait-freedom is a much stronger property which requires that all threads make progress in a finite number of steps.
@@ 2.3 Overview of Data Structures
In Table 1, we characterize common data structures based on their speed, (original) compatibility with SMR schemes, and compatibility when using our method. We omit hash maps as they are mostly arrays of Harris’ or Harris-Michael lists.
Table 1: Compatibility and performance characteristics of common non-blocking data structures.
Data Structure Performance EBR Original HP/IBR/HE HP/IBR/HE with SCOT
@@
These data structures employ “logical” deletion, i.e., a node inside a data structures is first simply marked (by stealing one bit from a node pointer), and then unlinked from the data structures (i.e., “physically” deleted). The original Harris’ List and Fraser’s Skip List use optimistic traversals, i.e., a chain of consecutive logically deleted nodes is deleted using one CAS operation; they allow implementing read-only search operations, which simply skip logically deleted nodes when iterating through the data structure. These optimizations improve performance and reduce contention.
@@
As we show in the table, data structures with optimistic traversals are incompatible with HP, IBR, and HE, a problem that we address in this paper.
@@ 2.4 Harris’ and Harris-Michael Linked Lists
5        node=malloc(sizeof(node_t));
6        node->Next = nullptr;
7        Head = node;
@@
23        node_t *curr, *next;
@@
28 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
@@
43   if ( !is_search && prev_next != curr ) // Skip for Search
@@
47            smr_retire(prev_next);
@@
In Search, the underlying Do_Find operation skips L41. In Insert and Delete, we also need to locate the right position for the key. Moreover, Insert and Delete ensure that we are not modifying pointers nodes which belong to nodes which are already (logically) deleted by cleaning up the discovered chain of logically deleted nodes in L41-L47 first. If we simply integrate HP without any modifications, L44 may crash while retrieving the next pointer in a loop.
3 Natarajan-Mittal Tree
@@ 4.2 SCOT for Harris’ List
The HP reclamation provides protect to safely retrieve the object and make a corresponding reservation. Additionally, we implement duplicate to duplicate the existing hazard pointer when moving to the next iteration. It is crucial to duplicate hazard pointers such that the old HP index has a lower numerical value than the new HP index (e.g., Hp0 to Hp1). This allows to avoid a small race window when iterating in the same (ascending) order of indices in retire.
Once we get to the dangerous zone of logically deleted nodes, Figure 4 runs a check in L16 which will ensure that reservation made in L8 is correct. Once the chain (or its subchain) gets unlinked, L16 fails and we restart the operation from the very beginning. We highlight in yellow color all major changes related to SCOT in HP.
1 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
@@
7        if ( curr == nullptr )  break ;
@@
9        if ( !isMarked(next) )
10            if ( curr->Key $\geq$ key )  break ;
11            prev = &curr->Next;
12            prev_next = next;
            hp.duplicate(Hp1, Hp2);
             // Duplicate Hp1 (curr) to Hp2
              hp.duplicate(Hp0, Hp3);
             // Duplicate Hp0 (next) to Hp3
13
14       else
            // Dangerous zone (marked): Validate that last safe node (N1) still points to the first unsafe (N2) node
@@
16
17       curr = getUnmarked(next);
        hp.duplicate(Hp0, Hp1);
         // Duplicate Hp0 (next) to Hp1
18
19
@@
19
20    if ( !is_search && prev_next != curr ) // Skip for Search
21        if ( !CAS(&prev, prev_next, curr )  goto …;
22        ;
@@
24            node_t *node = getUnmarked(prev_next->next);
@@ 4.3 SCOT for Natarajan-Mittal Tree
Due to complexity of the tree algorithm and very similar checks for optimistic traversals, we omit full pseudocode changes in this paper.

### v2

@@ Abstract
We present a new technique, Safe Concurrent Optimistic Traversals (SCOT), to address a well-known problem related to optimistic traversals with both classical and more recent memory reclamation schemes, such as Hazard Pointers (HP), Hazard Eras (HE), Interval-Based Reclamation (IBR), and Hyaline. Unlike Epoch-Based Reclamation (EBR), these schemes guarantee protection against stalled threads (robustness) but lack support for well-known data structures with optimistic traversals such as Harris’ original list, Natarajan-Mittal tree, among others. For these reclamation schemes, existing data structure implementations are either buggy (e.g., Natarajan-Mittal tree) or come with performance trade-offs (e.g., Harris-Michael modified list).
A recent work, HP++, supports optimistic traversals but uses a different API and is generally slower than even HP, not to mention more recent schemes such as IBR or Hyaline. Moreover, it has undesirable applicability trade-offs, and more complex implementation, among other issues. We propose a different method which keeps existing reclamation schemes intact but instead relies on data structure adaptations. Unlike existing Harris-Michael approach or HP++, our method retains performance benefits of the original data structure and also does not compromise performance of the underlying reclamation scheme. In fact, for IBR and Hyaline, our results almost match those of EBR, which often serves as a practical upper bound due to its great performance.
We implement and evaluate two fundamentally different data structures: Harris’ list and Natarajan-Mittal tree. SCOT enables their first correct implementations with optimistic traversals for HP, HE, IBR, and Hyaline. Our method is also directly applicable to other data structures, e.g., Fraser’s skip list or hash map that uses Harris’ list internally. We observe similar performance speedups (e.g., Harris vs. Harris-Michael list) that were previously available only to EBR users. As we discuss in the paper, our version of Natarajan-Mittal tree is the first bug-free implementation for HP, HE, IBR, and Hyaline; it achieves very high throughput, comparable to that of EBR.
@@ 1 Introduction
When using manual memory management, memory has to be safely reclaimed by freeing memory only when all ongoing operations with stale pointers complete. A number of past researchers proposed various safe memory reclamation (SMR) techniques in C, C++, or Rust. Non-blocking SMR progress is often crucial for the overall progress guarantees, as we do not want to compromise data structure progress properties. Unfortunately, many SMR schemes are trying to attain several highly desirable properties simultaneously while often unknowingly compromising non-blocking progress or other important properties one way or the other. A recent work [22] sheds more light onto this conundrum: it proves that it is impossible to achieve three highly desirable SMR properties simultaneously, i.e., only two properties can ever be achieved: (A) robustness (bounded memory usage), (B) easy integration (no dependency on any special APIs, such as roll-back mechanisms or OS-based primitives), and (C) wide applicability (compatibility with most non-blocking data structures).
@@
While striving to have (A) for non-blocking progress, [^1] If memory usage is unbounded, no further system progress can be made once memory is exhausted, regardless of whether actual locks are present. past researchers would compromise either (B), (C), or other properties that are tangentially related to non-blocking progress, sometimes without fully realizing that. To make matters worse, many researchers prioritize (C) because, intuitively, their goal is to support as many data structures as possible. However, we argue that sacrificing (C) rather than (B) by carefully adapting existing problematic data structures is a “necessary evil”, which is a crucial new insight. First, (C) in its strongest form – applying to all data structures – only currently known to exist in blocking SMR schemes [22], which is a non-starter for us. Second, even if we accept (C) in its relaxed form – applying to most data structures – the absence of (B) would lead to either inability to achieve non-blocking progress in practice due to dependence on OS-based primitives, lack of universally-defined APIs, or both. (We discuss some specific examples in Section 6.) Third, the absence of (C) simply implies that adaptations are required, not impossibility of the implementation.
@@
EBR [9] is a fast and easy-to-use approach but has a serious drawback that any stalled thread results in unbounded memory usage, i.e., it lacks (A). Once exhausting memory, the system must either reboot or block until memory is freed. Many researchers prefer EBR’s simplicity but attempt to solve its robustness issue. Unfortunately, such approaches typically sacrifice strict non-blocking progress [5, 23, 4, 1, 2], as discussed in Section 6, or need non-trivial roll-back mechanisms [21], which are fairly hard to generalize in terms of the API within the same library framework.
@@
HP [14] is an antipode of EBR: it is somewhat slower and more difficult to use but provides robustness with strict non-blocking guarantees by placing a hard limit on the number of local pointers within each thread. Prior to this work, well-known data structures such as Harris’ linked list [10] with optimistic traversals [^2] The original paper [10] did not explicitly mention that search can be done in a read-only manner (w/o any CAS operations). In this paper, by Harris’ list, we denote the enhanced version [11] with optimistic read-only traversals. and Natarajan-Mittal’s binary search tree (BST) [16] were not known to work [^3] Despite presenting Natarajan-Mittal Tree results in [26, 18] for HP and other SMR schemes, it was later argued [3] that such implementations are buggy and leak memory. Moreover, [3] pointed out two bugs, and only one bug was resolved in [3]. Prior to our work, the second bug related to optimistic traversals remained unresolved. with HP [12, 3], which is due to HP’s lack of the wide-applicability (C) property. A number of other recent schemes, including Interval Based Reclamation (IBR) [26], Hyaline-1S [18], and Hazard Eras (HE) [20], sometimes offer better performance but have similar issues to HP. HP++ [12] is another recent scheme. HP++ is slower than HP and, strictly speaking, makes compromises with both (B) and (C). However, HP++ still demonstrates support for these two data structures. We are inspired by HP++’s success with these data structures, which is indicative that many existing data structures can still be implemented, with an additional effort, despite prior beliefs regarding limited applicability of the SMR schemes in this category.
@@
This paper makes the following contributions. First, we depart from the typical strategy of designing a “silver-bullet” SMR approach for the reasons described above. Instead, we argue that the existing problematic data structures can be modified to accommodate SMR schemes with strict non-blocking properties. We demonstrate our new method in Section 3 for Harris’ list, Natarajan-Mittal tree, and other data structures. In Section 5, we evaluate these data structures with HP, HE, IBR and Hyaline-1S. We also compare results to EBR and show that performance benefits can be preserved in the same manner. Finally, in Section 6, we also touch upon how limitations of our approach with respect to wait-free traversals can be addressed.
@@ 2.1 Lock- and Wait-Freedom
The most widely used non-blocking progress is lock-freedom, which allows threads to interfere with each other while still guaranteeing that at least one thread makes progress in a finite number of steps. (It should be noted that good lock-free algorithms allow multiple threads to make concurrent progress in practice.) This should not be confused with “lockless” approaches that simply avoid explicit locks, where one preempted thread can block all other threads. Wait-freedom is a much stronger property which requires that all threads make progress in a finite number of steps.
@@ 2.3 Overview of Data Structures
In Table 1, we characterize common data structures based on their speed, (original) compatibility with SMR schemes, and compatibility when using our method. We omit hash maps as they are simply arrays of Harris’ or Harris-Michael lists.
Table 1: Subjective performance characterization of common non-blocking data structures and their compatibility with SMR schemes.
Data Structure Fast EBR HP/HE/IBR/Hyaline-1S HP/HE/IBR/Hyaline-1S with SCOT
@@
These data structures employ “logical” deletion, i.e., a node inside a data structure is first simply marked (by stealing one bit from a node pointer), and then unlinked from the data structures (i.e., “physically” deleted). The original Harris’ List and Fraser’s Skip List use optimistic traversals, i.e., a chain of consecutive logically deleted nodes is deleted using one CAS operation; they allow implementing read-only search operations, which simply skip logically deleted nodes when iterating through the data structure. These optimizations improve performance and reduce contention.
@@
As we show in the table, data structures with optimistic traversals are incompatible with HP, HE, IBR, and Hyaline-1S, a problem that we address in this paper.
@@ 2.4 Harris’ and Harris-Michael Linked Lists
5        Head = malloc(sizeof(node_t));
6        Head->Next = nullptr;
7        Head->Key = $\infty$;
@@
23        node_t **prev, *curr, *next;
@@
28 bool Do_Find(key_t key, node_t ***p_prev, node_t **p_curr, node_t **p_next, bool srch)
@@
43   if ( !srch && prev_next != curr ) // Search: skip
@@
47            smr.retire(prev_next);
@@
In Search, the underlying Do_Find operation skips L41. In Insert and Delete, Do_Find locates the right position for the key. Moreover, Insert and Delete ensure that we are not modifying the pointers of nodes that have been (logically) deleted by cleaning up the identified chain of logically deleted nodes in L41-L47 first. If we simply integrate HP without any changes, L44 may crash while retrieving the next pointer in a loop.
2.5 Natarajan-Mittal Tree
@@ 3.2 SCOT for Harris’ List
The HP reclamation provides a protect function to safely retrieve an object and make the corresponding reservation. Additionally, we implement dup to duplicate the existing hazard pointer when moving to the next iteration. It is crucial to duplicate hazard pointers such that the old HP index has a lower numerical value than the new HP index (e.g., Hp0 to Hp1). This allows to avoid a small race window when iterating in the same (ascending) order of indices in retire. (An alternative to dup would be index renaming via an indirection array. From our empirical observations, we found duplication to be generally cheaper.)
One tricky part is to avoid a premature duplication into Hp3, which would induce a memory barrier not present in Harris-Michael approach, potentially making the algorithm more costly. We achieve that by unrolling the loop in the original Harris’ algorithm and splitting it into two phases: (1) iteration through the safe zone and (2) iteration through the dangerous zone. Phase 1 only duplicates (shifts) prev and curr hazard pointers from curr and next, correspondingly. This is analogous to Harris-Michael algorithm. Upon leaving Phase 1, curr is duplicated (Hp3) in L18. After going to Phase 2, we no longer duplicate curr into prev (Hp2), which give us an extra benefit compared to Harris-Michael approach as the number of memory barriers through the dangerous zone is reduced.
1 bool Do_Find(key_t key, node_t ***p_prev, node_t **p_curr, node_t **p_next, bool srch)
@@
8        do
9            if ( curr && curr->Key $\geq$ key )  goto 4 ;
10            prev_next = nullptr;
11            prev = &curr->Next;
            hp.dup(Hp1, Hp2);
             // Duplicate Hp1[curr] to Hp2
12            curr = getUnmarked(next);
13            if ( curr == nullptr )  goto 4 ;
            hp.dup(Hp0, Hp1);
             // Duplicate Hp0[next] to Hp1
@@
15
16       while !isMarked(next);
         prev_next = hp.dup(Hp1,Hp3);
         // Hp1[curr] to Hp3
        // Dangerous zone (marked): Check last safe node (N1) still points to the first unsafe (N2) node
17          do
18              curr = getUnmarked(next);
19              if ( curr == nullptr )  goto 4 ;
              hp.dup(Hp0, Hp1);
             // Duplicate Hp0[next] to Hp1
20              next = hp.protect(curr->Next, Hp0);
@@
22
23       while isMarked(next);
24
@@
26
    // Skip for search
27    if ( !srch && prev_next != nullptr && prev_next != curr )
28        if ( !CAS(&prev, prev_next, curr )
29            goto 4;
30
@@
32            node_t node;
33            node = getUnmarked(prev_next->next);
@@
Once we get to the dangerous zone of logically deleted nodes, Figure 4 runs a check in L24 which will ensure that reservation made in L23 is correct. Once the chain (or its subchain) gets unlinked, L24 fails and we restart the operation from the very beginning. We highlight in the green color all major changes related to SCOT in HP. We use pink to highlight the dangerous zone traversal.
3.2.1 Recovery Optimization
For simplicity, we have stated that when validation fails, we need to go back to the beginning of the list. While this is acceptable, this validation check is not always critical: the last safe node may simply point to another node now (e.g., when a new node was inserted or the chain of logically deleted nodes was already eliminated by a concurrent thread). In such cases, we can simply escape from the dangerous zone and continue to new node. We must, however, ensure that the last safe node is still not logically deleted. When it is deleted, the last safe node (prev) is in a dangerous zone itself (though still safe to access due to prior reservation). Depending on the SMR algorithm, we have different choices.
For HP and HE, protect uses indices and cancels previous reservations. As keeping a predecessor of prev would induce extra hazard pointers and duplication (hence memory barriers), it makes sense to simply go back to the beginning of the list when prev gets logically deleted.
For IBR and Hyaline-1S, protect has a cumulative effect and does not cancel previous reservations. For these schemes, we implement a small on-stack ring buffer of prev pointers, which allows us to fall further back in the list when prev gets logically deleted. Empirically, we found that a ring buffer of size 8 is more or less optimal. Only when exhausting all 8 predecessors, do we go back to the beginning of the list. (See Figure 6.)
@@ 3.2.1 Recovery Optimization
Figure 6: Recovery Optimization: IBR and Hyaline-1S can go from N1 all the way back to N”.
@@ 3.3 SCOT for Natarajan-Mittal Tree
Unlike Harris’ list, empirically, we found that recovery does not help improve performance for various key ranges, primarily because of the hierarchical structure of the tree (e.g., when there is a mismatch, likely, the tree has diverged substantially anyway). If ancestor no longer points to the original successor, we simply restart the search from the very beginning.
Due to complexity of the algorithm and very similar checks for optimistic traversals, we omit pseudocode in this paper. In Section 5, we evaluate Natarajan-Mitttal Tree with SCOT.